### v1

# Solid-State Dewetting of Polycrystalline Thin Films: a Phase Field Approach

Abstract
Solid-state dewetting (SSD) is the process by which thin solid films break up and retract on a substrate, resulting in the formation of nanostructures and islands. While SSD in single-crystalline films is generally understood as a surface-energy-driven process mediated by surface diffusion, polycrystalline films feature additional complexity due to the presence of grain boundaries. Theoretical investigations and simulation frameworks have mainly focused on single-crystalline SSD. In this work, we present and apply a multi-phase-field framework that captures key mechanisms in polycrystalline thin films with multiple grains and grain boundaries in three dimensions. By considering isotropic surface and interface energy and curvature-driven grain boundary motion, we demonstrate how the framework reproduces the key phenomenology of SSD while being consistent with predictions based on energetic arguments. We also introduce refined analytical criteria for the onset of dewetting, confirmed by numerical simulations. Key evidence is given concerning the onset of dewetting at grain-boundary triple junctions. Moreover, we showcase SSD of polycrystalline patches. This work paves the way for in-depth investigations of the morphological evolution of polycrystalline thin films.
Thin crystalline films are metastable due to their high surface-to-volume ratio. At high temperatures below melting, they typically break up and form three-dimensional agglomerates in a process known as solid-state dewetting (SSD) [1]. While detrimental to planar architectures, this phenomenon realizes a unique kinetic pathway for the self-assembly of complex nanostructures. Various strategies have been proposed to control the outcomes of SSD in single-crystalline thin films, such as patterning the film or the underlying substrate [2, 3, 4, 5]. In these systems, the dynamics are predominantly governed by surface diffusion [1].
Several technologically relevant thin films are polycrystalline [6, 7]: they consist of crystalline domains (grains) with varying crystallographic orientations, which are separated by interfaces known as grain boundaries (GBs). Their morphological evolution is influenced not only by surface diffusion, as in single crystals, but also by additional effects associated with GBs [8, 9, 10, 11]. Arguably, the most relevant one is that surface grooves develop at GBs. Moreover, GBs migrate through a mechanism resembling attachment-detachment kinetics, with a velocity governed largely by capillarity (mean curvature flow) and potentially influenced by elasticity effects [12, 13].
Modeling and numerical simulation played a critical role in understanding and tailoring single-crystalline SSD (see, e.g., [14, 4, 5]). Studies on GB grooving and few-grains microstructures have also yielded valuable insights [15, 16, 17, 18, 19], but remain limited in capturing the coupled dynamics of GB migration and surface diffusion across many grains and long timescales. Phase field (PF) models [20, 21, 22, 23] hold a prominent position to study SSD. They realize an implicit description of geometries via smooth order parameter(s), which naturally deals with topological changes, complex geometries, and three-dimensionality, all key aspects of SSD [24, 4, 5, 25, 26]. Convenient PF models coupling diffusion, phase transformation, and interface migration have been proposed [27, 28, 18, 29, 30]. A recent extension of the grand-potential PF model [27, 28, 31] accurately captures bulk and surface diffusion coupled with interface migration, reproducing key features of polycrystalline SSD, such as thermal grooving at GBs [29].
In this work, we present a multi-PF (MPF) approach to studying SSD of polycrystalline thin films. After a brief overview of the model’s key features, we present its capabilities and validate it through benchmarks. Furthermore, we discuss novel insights from simulations and semi-analytic equations regarding the onset of SSD. We also apply the MPF model to prototype polycrystalline films and patches, outlining differences and similarities with their single-crystalline counterparts.
We employ the MPF model introduced in [29], considering its specialization to systems dominated by surface/interface diffusion, as well as interface migration, while neglecting bulk diffusion. We also consider grains of the same pure material and neglect any elasticity effects that may arise from GB migration. In brief, order parameters $\{\phi_{1}(\mathbf{r}),\dots,\phi_{N}(\mathbf{r})\}\equiv\boldsymbol{\varphi}$ akin to smooth characteristic functions track different domains/phases. To account for material transport, global concentration fields for $K$ elements $c_{i}(\mathbf{r})$ are considered. They can be related to the composition of each phase ($c_{i\alpha}$) via interpolation $c_{i}(\mathbf{r})=\sum_{\alpha=1}^{N}\phi_{\alpha}(\mathbf{r})c_{i\alpha}$ (spatial dependence will be omitted for brevity). In order to describe the chemical interactions within the MPF system, we define a set of $K-1$ independent chemical potential fields, $\{\mu_{1},\dots,\mu_{K-1}\}\equiv\boldsymbol{\mu}$, corresponding to the concentration fields, which distinguish between the solid film, the substrate, and the surrounding vacuum. The grand potential functional reads
$\begin{split}\Psi[\boldsymbol{\varphi},\nabla\boldsymbol{\varphi},\boldsymbol{\mu}]=&\int_{\Omega}\bigg{[}f_{\rm I}(\boldsymbol{\varphi},\nabla\boldsymbol{\varphi})+\psi(\boldsymbol{\varphi},\boldsymbol{\mu})\bigg{]}\mathrm{d}\mathbf{r},\\ \end{split}$ (1)
where
$f_{\rm I}=\sum_{\alpha}^{N}\sum_{\beta=\alpha+1}^{N}\gamma_{\alpha\beta}\left(\epsilon\nabla\phi_{\alpha}\cdot\nabla\phi_{\beta}+\frac{16\mathcal{G}(\boldsymbol{\varphi})\phi_{\alpha}\phi_{\beta}}{\pi^{2}\epsilon}\right),$ (2)
encodes the interfacial energy with $\epsilon$ the interface thickness, $\mathcal{G}(\boldsymbol{\varphi})=1$ if $\sum_{\alpha}\phi_{\alpha}=1$ and $\phi_{\alpha}\geq 0,\ \forall\alpha$, while $\mathcal{G}(\boldsymbol{\varphi})=\infty$ otherwise (realizing a double-obstacle potential). The function $\psi(\boldsymbol{\varphi},\boldsymbol{\mu})$ corresponds to the grand-potential density
$\psi(\boldsymbol{\varphi},\boldsymbol{\mu})=\sum_{i=1}^{K-1}\Big{[}\sum_{\alpha=1}^{N}(f_{i\alpha}-\mu_{i}c_{i\alpha})\phi_{\alpha}\Big{]}.$ (3)
In this study we used a parabolic free energy density $f_{i\alpha}=A_{i}^{\rm chem}(c_{i\alpha}-c_{i\alpha}^{\text{eq}})^{2}$ depending on the equilibrium concentrations $c_{i\alpha}^{\text{eq}}$ and constants $A_{i}^{\rm chem}$.
Interface diffusion can be described by $\partial_{t}c_{i}=-\nabla\cdot\mathbf{J}_{i}$ and $\mathbf{J}_{i}$ material flux for the $i$-th element constrained at the interfaces. Upon assuming quasi-equilibrium, a uniform chemical potential for the $i$-th species across different phases is considered, namely $\mu_{i}=\frac{\partial f_{i1}}{\partial c_{i1}}=\ldots=\frac{\partial f_{iN}}{\partial c_{iN}}$. A Legendre transformation can be performed to substitute the chemical potential as a fundamental variable governing time evolution $c_{i\alpha}(t)=\frac{\mu_{i}(t)}{2A_{i}}+c_{i\alpha}^{\text{eq}}$. Phase boundary migration is accounted for through the dynamics of $\phi_{\alpha}$, as in classical PF models, with $\Psi$ being the energy functional. The resulting model equations read
$\begin{split}\epsilon\partial_{t}{\phi}_{\alpha}=&\frac{1}{\tilde{N}}\sum_{\beta=1}^{N}m_{\alpha\beta}\left(\frac{\delta\Psi}{\delta\phi_{\beta}}-\frac{\delta\Psi}{\delta\phi_{\alpha}}\right),\\ \frac{1}{2A_{i}}\partial_{t}\mu_{i}=&\nabla\bigg{[}\sum_{j=1}^{K-1}\sum_{\alpha=1}^{N}\sum_{\beta=\alpha+1}^{N}\frac{32}{\epsilon\pi^{2}}M_{\alpha\beta j}\phi_{\alpha}\phi_{\beta}\nabla\mu_{j}\bigg{]}\\ &-\sum_{\alpha=1}^{N}c_{i\alpha}^{\text{eq}}\partial_{t}\phi_{\alpha},\end{split}$ (4)
with $m_{\alpha\beta}$ and $M_{\alpha\beta}$ the mobility and diffusivity at the $\alpha\beta$-interface, respectively. The timescale will be expressed in arbitrary units. Sharp interface limits showing convergence to the targeted kinetics (surface diffusion coupled with attachment-detachment kinetics) can be found in [29, 32]. Simulations are obtained by numerically solving Eqs. (4) using a finite difference integration scheme, implemented in the modular software package for multi-physics, parallel computations PACE3D [33]. Periodic boundary conditions are applied in the in-plane directions, while no-flux Neumann conditions are imposed at the top and bottom boundaries.
Figure 1: Modeling and 2D simulation of polycrystalline SSD. (a) Scheme of the considered geometries and labeling of the phases as considered in the MPF model. The grain aspect ratio is defined as $r^{\rm 2D}=W/H$. (b) SSD of a thin film with three grains. The central one has an aspect ratio of $r^{\rm 2D}=25$. The two at the border of the film have both $r^{\rm 2D}=12.5$. The timescale is relative to the one needed to reach equilibrium. (c) MPF simulations of SSD with varying $r^{\rm 2D}$ for a film ideally formed by grains of the same size. A critical aspect ratio of $r^{\rm 2D}_{\rm c}\approx 9$ is obtained in simulations.
Figure 2: Geometrical parametrization and estimation of the critical aspect ratios. (a) Schematics of 2D equilibrium morphology. The dashed line illustrates the thickness ($H$) of the conformal initial film. (b) Schematics (top view) of the idealized setting of hexagonal grains in 3D with equal size. (c) Spherical caps truncated by $n$-gons. (d) $r_{\rm c}(\theta)$ in 2D and 3D as predicted by Eq. (5) and Eq. (8) for $d=0$ (sharp interface limit) and $d>0$ as realized in the MPF simulations. The vertical dotted line indicates the $\theta$ realized in the MPF simulations. Symbols show $r_{\rm c}$ obtained in 2D (circle) and 3D (square) MPF simulations.
In our polycrystalline thin film model, the order parameters $\phi_{\alpha}$ track the substrate, the vapor phases, and the different grains within the film. Fig. 1a shows an illustration of a polycrystalline film composed of three grains, where we label the vacuum phase $\alpha=1$, the substrate $\alpha=N$, and the grains $1<\alpha<N$. We refer to the ratio of width to height $r^{\rm 2D}=W/H$ as the aspect ratio of the 2D grains. We use a two-concentration-field model for the substrate and the film with $c_{\text{substrate},N}^{\text{eq}}=1$ and $c_{\text{grain},\alpha}^{\text{eq}}=1$. We set $A_{\text{substrate}}^{\rm chem}=50$ and $A_{\text{grain}}^{\rm chem}=5$ [32]. Interfacial (surface) diffusion is only allowed at the film-vapor interface via $M_{\alpha\beta}=\delta_{1\beta}$. GB mobilities are set to $10^{-3}$ while all other mobilities set to $1$ and $\gamma_{\alpha\beta}=0.5$ for all interfaces.
A first representative simulation of SSD for a 2D three-grains film, with aspect ratios $r^{\rm 2D}_{\text{left}}=r_{\text{right}}^{\rm 2D}=12.5$ and $r_{\text{center}}^{\rm 2D}=25$, is shown in Fig. 1b. This setting approximates stripes infinitely extended in the direction perpendicular to the illustration [5]. In the initial stages, rims form at the edges of the film as it retracts, while a groove develops at the GBs where the film eventually breaks. Afterward, every grain evolves following the SSD dynamic of a single crystal, resulting in a number of final islands equal to the number of initial grains. In the central grain at $t=0.25$, the typical material depletion in the center due to surface diffusion [1] can be observed, leading to a dent in the surface. For grains with large aspect ratios, as obtained after recrystallization, impingement of single-crystal patches, or abnormal grain growth, material depletion can cause breakpoints within grain centers, leading to more islands than initial grains.
Figure 3: Simulations (3D) of morphological evolution and SSD in polycrystalline thin films with an idealized initial microstructure. (a) Perspective view of the equilibrium configuration for thin films with varying aspect ratio, except for $r^{\rm 3D}=16$ for which an intermediate stage is reported. (b) Representative stages of SSD for grains with $r^{\rm 3D}=8$ (top view). The timescale is expressed relative to the breakup time. (c) Representative stages of SSD for grains with $r^{\rm 3D}=32$ (top view). The timescale is set as in panel (b).
The well-known Young’s law can be applied to calculate the contact angles at the GB. There, the grooves exhibit an (equilibrium) angle $\theta$ w.r.t the planar surface of the initial film that satisfies the identity $2\gamma_{\rm s}\sin\theta=\gamma_{\rm b}$ where $\gamma_{\rm s}$ and $\gamma_{\rm b}$ are the (isotropic) surface energy density of the crystal surface and the boundary, respectively. With the chosen interfacial energies ($\gamma_{\rm s}=\gamma_{\rm b}=0.5$), it is $\theta=\pi/6$ and is well reproduced by the MPF simulations. For a GB between two infinitely extended grains ($r^{\rm 2D}\rightarrow\infty$), the depth of the grooves monotonously increases over time [8], thus inherently leading to breakup. For a finite aspect ratio, however, surface diffusion and thus grooving at the GBs stops when the equilibrium shape of the surface of the grain(s) is reached [10]. For isotropic surface energy, the equilibrium shape corresponds to a circular profile that satisfies the equilibrium angle $\theta$ at the groove. By simulating a periodic system with uniformly distributed grains and systematically increasing their aspect ratio, we identify a critical aspect ratio $r_{\rm c}^{\rm 2D}$ above which SSD occurs ($r^{\rm 2D}>r_{\rm c}^{\rm 2D}$). It results approximately $r_{\rm c}\approx 9$ for the chosen parameters, as illustrated in Fig. 1c. We remark that this occurs when the (diffuse) interface meets the substrate.
Following Ref. [10], $r_{\rm c}^{\rm 2D}$ can be independently estimated to validate the simulation results and the underlying framework. For an interface of infinitesimal thickness, $r_{\rm c}^{\rm 2D}$ is determined by considering mass conservation, i.e., equating the initial grain area to the area of a circular cap with contact angle $\theta$ and base $W$, i.e. $A_{\rm G}=A_{\rm C}$, with $A_{\rm G}=WH$ and $A_{\rm C}=W^{2}(2\sin\theta)^{-2}(\theta-\sin\theta\cos\theta)$. For an interface with a finite effective thickness, where breakup occurs when the groove reaches a distance $D$ from the substrate, an additional area $A_{D}=WD$ must be accounted for. The mass conservation condition then reads $A_{\rm G}=A_{\rm C}+A_{\rm D}$. Expressing lengths as fractions of the initial film thickness ($W=rH$ and $D=dH$, implying $d<1$) and solving for the (critical) aspect ratio, we get
$r_{\rm c}^{\rm 2D}=\frac{4\sin^{2}\theta}{\theta-\sin\theta\cos\theta}(1-d).$ (5)
For $d=0$ and $\theta=\pi/6$ this results into $r_{\rm c}\approx 11$. By setting $d=\epsilon/H=0.2$, we get $r_{\rm c}\approx 8.8$, which is in excellent agreement with the simulation results. This agreement also supports the use of Eq. (5) to predict the sharp interface limit, $d=0$.
We extend here the previously outlined geometrical and energetic arguments to three dimensions. In an idealized scenario where regular $n$-gon shaped grains of uniform size perfectly tile the substrate [10, 30] (see Fig. 2b for $n=6$) the aspect ratio can be defined as $r^{\rm 3D}=W/H$ where $W=2a$ and $a$ is the apothem of the $n$-gon. The volume of each grain with height $H$ is then given by $V_{\rm G}=A_{n}H$ with the base area $A_{n}=n\tan(\pi/n)W^{2}/4$. As an ideal equilibrium condition, the system is assumed to evolve toward a spherical cap truncated by the $n$-gon (see Fig. 2c). As derived in the Supplementary Material (SM), the corresponding integral expression for its volume is
$V_{\rm C}=\frac{nW^{3}}{4}\int_{0}^{\tilde{R}}\tilde{z}_{n}(\theta,\tilde{\rho})\vartheta_{n}(\tilde{\rho})\tilde{\rho}\,\mathrm{d}\tilde{\rho},\$ (6)
with
$\tilde{z}_{n}(\theta,\tilde{\rho})=\sqrt{\sin^{-2}(\theta)-\tilde{\rho}^{2}}-\sqrt{\sin^{-2}(\theta)-\tilde{R}^{2}},$ (7)
$\tilde{\rho}=\rho/a$, $\tilde{R}=R/a=1/\cos(\pi/n)$, $\vartheta_{n}(\tilde{\rho})=\pi/n$ if $\tilde{\rho}<1$ and $\pi/n-\arccos(\tilde{\rho}^{-1})$ otherwise. By imposing mass conservation, $V_{\rm G}=V_{\rm C}+V_{\rm D}$ with $V_{\rm D}=A_{n}D$ and $D=dH$ in analogy to Eq. (5), the critical aspect ratio results
$r_{\rm c}^{\rm 3D}=\frac{\tan(\pi/n)}{\tilde{V}_{\rm C}(n,\theta)}(1-d).$ (8)
Fig. 2d illustrates the dependence of the critical aspect ratios, Eqs. (5) and (8) for regular hexagonal grains ($n=6$), on $\theta$. Note that $r_{\rm c}^{\rm 3D}$ increases with $n$, indicating that grains with fewer edges undergo SSD earlier than those with more edges. This suggests that sharp corners promote film rupture. Note that from the equations above, $\tilde{z}=0$ is ensured at the junction where $\tilde{\rho}=\tilde{R}$. Since the (spherical) elevation decreases with radius, an enhanced grooving at junctions as observed in previous works [34, 30] is naturally accounted for. The elevation of the junction (J) and of the GB at the midpoint between junctions can be expressed as (see SM)
$\begin{split}d_{\rm J}&=H-z_{n}(\theta,R),\\ d_{\rm GB}&=\frac{W}{2}\left(\frac{1}{\tan{\theta_{c}}}-\sqrt{\frac{1}{\sin^{2}{\theta_{c}}}-\frac{1}{\cos^{2}({\pi}/{n})}}\right)+d_{\rm J}.\end{split}$ (9)
Figure 4: SSD of square polycrystalline patches with an arbitrary internal grain structure. The two panels (a) and (b) show representative stages of the SSD evolution with two different realizations of the initial conditions featuring different spreads in grain sizes and shapes but analogous qualitative behaviors. The timescales are arbitrarily expressed relative to the breakup time (similar in both the illustrated cases). Note that panel (b) illustrates the evolution until the very late stages.
MPF simulations of SSD of the idealized initial configurations as in Figs. 2b are presented in Fig. 3a varying $r^{\rm 3D}$. For relatively small aspect ratios, the morphology of grains at equilibrium is well described by spherical caps truncated by hexagons ($n=6$ in Fig. 2c). $d_{\rm J}$ and $d_{\rm GB}$ decrease with increasing $r^{\rm 3D}$ and correspond well to the prediction of Eq. (9). For instance, considering the case of $r^{\rm 3D}=5.5$, the predicted elevations normalised to the initial film thickness are $d_{\rm J}/H\approx 0.37$ and $d_{\rm GB}/H\approx 0.64$ and the simulated normalised elevations measured at the nominal interface between vapor phase and grains ($\phi_{1}=0.5$) are $d_{\rm J}/H\approx 0.37$ and $d_{\rm GB}/H\approx 0.65$. Moreover, a critical aspect ratio of $r^{\rm 3D}_{\rm c}\approx 6.7$ is obtained from simulations (c.f. Fig. 3a). Eq. (8), for the considered scenario where $\theta=\pi/6$, delivers an estimation of $r_{\rm c}^{\rm 3D}\approx 8.7$ for $d=0$ and $r_{\rm c}^{\rm 3D}\approx 6.8$ considering the diffuse interface of MPF 3D simulations with $d=\epsilon/H=0.22$ (see also Fig. 2d), in excellent agreement with the result of simulations. On the one hand, we can conclude that the MPF framework well reproduces estimation based on interface/surface energy minimization. On the other hand, predictions of Eqs. (8) and (9) are confirmed by (independent) numerical simulations. They can be readily exploited to provide initial estimates of the onset of SSD in polycrystalline thin films, given the aspect ratio of the grains. We note, however, that localized inhomogeneities and defects typically occur in polycrystals, which can trigger SSD but with a highly heterogeneous character.
Figure 3b shows a representative morphological evolution during dewetting. It is obtained by a MPF simulation with the idealized initial configuration shown in Fig. 2b. A relatively large $r^{\rm 3D}$ is chosen to ensure film rupture. SSD initiates at junctions, confirming seminal equilibrium considerations [10], other recent simulations [30], and predictions discussed above. Additionally, a detailed account of the dynamics at GB junctions is obtained, accurately capturing morphological evolution, including the formation and growth of non-trivial opening holes. At intermediate stages, as in 2D, grains detach from neighbors and evolve into single-crystalline patches. A similar qualitative behavior, enriched by additional features, emerges at larger aspect ratios; see Fig. 3c. Breakup initiates at triple junctions and rapidly follows at grain boundaries. When the grains detach completely from each other, the dynamics converge to the one expected for a single crystalline patch. A rim forms at the edges (see also Fig. 3a for $r^{\rm 3D}=16$). For the considered aspect ratio, a breakup at the center of the patch is obtained, resembling a classical feature of SSD of single-crystalline patches  [2, 24, 4]. A semi-toroidal shape forms and, driven by curvature gradients, evolves until the central hole eventually closes.
We finally present the case of a polycrystalline thin film of finite size (a patch); see Fig. 4, which reports the SSD dynamics for two different initial conditions. The grains are initialized with varying morphologies and aspect ratios. As observed earlier, rims form at the grain boundaries, while holes open at junctions. Such holes initiate dewetting fronts within the patch, while a continuous rim effectively develops along the patch’s edges, breaking only at later stages. Aside from the rapid disappearance of small grains, the slower dynamics of rims and peripheral GBs, compared to hole opening, are evident along edges with both relatively small and large grains. This can be seen, for instance, by comparing the upper and lower edges in Fig. 4a, and it is also obtained for a more heterogeneous distribution of grains, as shown in Fig. 4b. We thus find that SSD in polycrystalline patches leads to early-stage morphologies characterized by material accumulation between the center and the initial patch boundaries, closely resembling the behavior observed in single crystals [24, 4]. The presence of GBs eventually leads to uncontrolled breaking within the patch for usual polycrystals, where grain contributions are dominated by randomness. However, the design of GB networks may lead to additional degrees of freedom to drive self-assembly in a controlled manner. In this context, we remark that the opening of holes occurs relatively fast compared to the rim retraction mediated by surface diffusion (see, e.g., Fig. 4b).
In conclusion, we demonstrated a MPF tackling the complexity of SSD dynamics in polycrystalline thin films. The simulation results are consistent with independent predictions based on energetic arguments and successfully reproduce the key phenomenology of surface diffusion and curvature-driven GB motion in systems with multiple grains up to equilibrium. We further elucidated key aspects of the onset of dewetting and analyzed the SSD of polycrystalline patches, highlighting the differences and similarities with single-crystalline SSD. Tailoring GB distributions can expand the design space for SSD-driven self-assembly. In this context, our approach offers a versatile tool to explore the associated parameter space. For the sake of generality, we have illustrated the framework using ideal, minimal settings, particularly concerning material parameters. However, by leveraging extended parameterizations and couplings already available in PF models, such as those addressing anisotropic surface or GB energy densities [35, 36, 37] as well as elasticity and plasticity effects [21, 38, 39], the proposed framework can be adapted to material-specific settings.
This work was funded by the German Research Foundation (DFG) under Project ID 390874152 (POLiS Cluster of Excellence). M.S. and A.V. acknowledge support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) within FOR3013, project number 417223351. D.S. and B.N. are grateful for support from the Helmholtz association within programme MSE (P3T1) number 43.31.01. The authors gratefully acknowledge the computing time made available to them on the high-performance computer at the NHR Center of TU Dresden. This center is jointly supported by the Federal Ministry of Education and Research and the state governments participating in the NHR.
Author Declarations
Conflict of Interest
The authors have no conflicts to disclose.
Author Contributions
Paul Hoffrogge: Data curation (equal), Investigation (equal), Methodology (equal), Software (equal), Visualization (equal), Writing – review & editing (equal). Nils Becker: Investigation (equal), Data curation (equal), Methodology (equal), Software (equal), Visualization (equal), Writing – review & editing (equal). Daniel Schneider: Software (support), Data curation (equal), Funding acquisition (equal), Methodology (support), Supervision (support). Britta Nestler: Software (support), Funding acquisition (equal), Supervision (lead) Axel Voigt: Conceptualization (support), Funding acquisition (equal). Marco Salvalaglio: Conceptualization (lead), Investigation (support), Funding acquisition (equal), Methodology (support), Supervision (support), Visualization (equal), Writing – original draft, Writing – review & editing (equal)
Data Availability
The data that support the findings of this article are openly available [40].
SUPPLEMENTARY MATERIAL
Solid-State Dewetting of Polycrystalline Thin Films: a Phase Field Approach
Paul Hoffrogge,1,2,∗, Nils Becker,3,∗, Daniel Schneider,1,2,3
Britta Nestler,1,2,3 Axel Voigt,4,5 and Marco Salvalaglio4,5,†
1Institute of Nanotechnology—Microstructure Simulations (INT-MS),
Karlsruhe Institute of Technology, Hermann-von-Helmholtz-Platz 1, 76344
2Institute for Digital Materials Science (IDM),
Karlsruhe University of Applied Sciences, Moltkestrasse 30, 76133 Karlsruhe, Germany
3Institute for Applied Materials (IAM-MMS),
Karlsruhe Institute of Technology, Strasse am Forum 7, 76131 Karlsruhe, Germany
4Institute of Scientific Computing, TU Dresden, 01062 Dresden, Germany
5Dresden Center for Computational Materials Science (DCMS), TU Dresden, 01062 Dresden, Germany
These authors contributed equally to this work
marco.salvalaglio@tu-dresden.de
Theory on N-gonal Polycrystalline Thin-Films
Angles and base area
(a) Side view
(b) Top view
Figure S-1: Spherical cross-section (a) and $n$-gon unit cell (b) of a grain in the tile pattern.
A grain with a spherical surface, having radius of curvature $R_{\kappa}$ as depicted in Fig. S-1a is assumed. The tangent angle $\theta_{t}$ is related to the radius of curvature as follows:
$r=R_{\kappa}\sin(\theta_{t}).$ (S-1)
Let us consider the grain to be truncated by a regular $n$-gon such that the grain is in contact with its $n$ neighbor grains at a distance, named the apothem $a$ of the $n$-gon. Here, we assume a constant contact angle $\theta$ is realized with all its neighbors, giving
$a=R_{\kappa}\sin(\theta),$ (S-2)
which then, using Eq. (S-1), implies
$\theta_{t}(r)=\arcsin\left(r\sin(\theta)/a\right).$ (S-3)
The outerscribed circle radius of the $n$-gon is denoted as $R$ (see Fig. S-1b). The area of the $n$-gons unit cell, is given by $al/2$, thus the total base area of the grain is $A_{n}=nal$ and since $l=a\tan(\pi/n)$, we have
$A_{n}=na^{2}\tan(\pi/n).$ (S-4)
This triangle also reveals the relation
$a$ $=R\cos(\pi/n).$ (S-5)
The vertices of the $n$-gon correspond to multi-junctions. The tangent angle of the surface at the junction is given by
$\theta_{\mathrm{junction}}=\theta_{t}(R)=\arcsin\left(\frac{\sin(\theta)}{\cos(\pi/n)}\right).$ (S-6)
The condition $\theta_{\mathrm{junction}}<\pi/2$ must be satisfied for the junction to exist; otherwise, immediate dewetting occurs. Therefore,
$\sin(\theta)<\cos(\pi/n)=\sin(\pi/2\pm\pi/n),$ (S-7)
is obtained, which gives the maximal contact angle
$\theta_{\mathrm{max}}=\pi/2-\pi/n\,.$ (S-8)
Volume of N-gonal grain
Let us assume $\theta<\theta_{\mathrm{max}}$, then the elevation of the surface $z_{n}$ of the $n$-gon is, according to Pythagorean theorem (see Fig. S-1a),
$(z_{n}(\theta,r)-z_{m}(\theta))^{2}+r^{2}=R_{\kappa}^{2}(\theta),$ (S-9)
and hence
$z_{n}(\theta,r)=z_{m}+\sqrt{R_{\kappa}^{2}(\theta)-r^{2}}.$ (S-10)
The elevation at which the junction pit forms is given as $z_{n}(\theta,R)$. Within the grain, the elevation is lowest at the junction, hence we enforce without loss of generality $z_{n}(\theta,R)=0$ to obtain elevations relative to the junction pit. This yields
$z_{m}=-a\sqrt{\sin^{-2}(\theta)-\cos^{-2}(\pi/n)}.$ (S-11)
The volume $V_{\rm C}$ of the $n$-gon grain featuring a spherical cap can be calculated by integration of the elevation in cylindrical coordinates:
$V_{\rm C}=2nV_{\text{unit cell}}=2n\int_{0}^{R}z_{n}(\theta,\rho)\vartheta_{n}(\rho)\rho\,\mathrm{d}\rho,$ (S-12)
where $\vartheta_{n}(\rho)$ is the angle between two edges of the $n$-gon unit cell (see Fig. S-1b),
$\vartheta_{n}(\rho)=\begin{cases}\pi/n&\rho<a\\ \pi/n-\arccos(a/\rho)&\rho\geq a\end{cases}\,.$ (S-13)
Introducing the dimensionless quantities $\tilde{z}_{n}=z_{n}/a$, $\tilde{R}={R}/a$, $\tilde{\rho}=\rho/a$ and $\tilde{V}_{\rm C}=V_{\rm C}/(2na^{3})$, we can rewrite
$V_{\rm C}=2na^{3}\tilde{V}_{\rm C}(\theta)=2na^{3}\int_{0}^{\tilde{R}}\tilde{z}_{n}(\theta,\tilde{\rho})\vartheta_{n}(\tilde{\rho})\tilde{\rho}\,\mathrm{d}\tilde{\rho},$ (S-14)
where
$\tilde{R}$ $=1/\cos(\pi/n),$ (S-15)
$\tilde{z}_{n}(\theta,\tilde{\rho})$ $=\sqrt{\sin^{-2}(\theta)-\tilde{\rho}^{2}}-\sqrt{\sin^{-2}(\theta)-\tilde{R}^{2}},$ (S-16)
$\vartheta_{n}(\tilde{\rho})$ =$=\begin{cases}\pi/n&\tilde{\rho}<1\\ \pi/n-\arccos(\tilde{\rho}^{-1})&\tilde{\rho}\geq 1\end{cases}\,.$ (S-17)
Critical aspect ratio
Let us consider an initially flat $n$-gonal grain with height $H$. Its volume is then given by
$V_{\rm G}=A_{n}H.$ (S-18)
Let us assume that the grain coarsens, eventually reaching the stable quasi-equilibrium truncated spherical shape (see also the main text, Fig. 2c). It is assumed that the film breaks up at some height $D<H$, giving the breakup volume
$V_{\text{breakup}}$ $=V_{\rm C}+A_{n}D.$ (S-19)
Therefore, the breakup is ensured at a finite time when the volume of the grains satisfies the following condition
$\begin{split}V_{\rm G}&<V_{\text{breakup}},\\ A_{n}H&<2na^{3}\tilde{V}_{\rm C}(\theta)+A_{n}D,\\ \frac{A_{n}}{2na^{3}}&<\frac{\tilde{V}_{\rm C}(\theta)}{H-D}.\end{split}$ (S-20)
Since the function $1/x$ is strictly montonously decreasing for $x>0$ it follows
$\frac{2na^{3}}{A_{n}}$ $>\frac{H-D}{\tilde{V}_{\rm C}(\theta)}.$ (S-21)
With $D$ expressed in terms of the film thickness, $D=dH$, with $d<1$ an adimensional factor, we have
$\frac{2na^{3}}{A_{n}H}>\frac{1-d}{\tilde{V}_{\rm C}(\theta)},$ (S-22)
and with $A_{n}=na^{2}\tan(\pi/n)$ we obtain
$\frac{2a}{H\tan(\pi/n)}>\frac{1-d}{\tilde{V}_{\rm C}(\theta)}.$ (S-23)
The critical aspect ratio $r^{\rm 3D}_{\rm c}$, defined as twice the apothem divided by the initial height $H$ of the film, $r^{\rm 3D}=2a/H$ (see also the main text), thus results
$r^{\rm 3D}_{\rm c}=\frac{2a}{H}=\frac{\tan(\pi/n)}{\tilde{V}_{\rm C}(\theta)}(1-d).$ (S-24)
Relative Depth of GB and Junction
The depth of the GB and junction w.r.t. the initial film height is given by $d_{\rm GB}=H-z_{n}(\theta,a)$ and $d_{\rm J}=H-z_{n}(\theta,R)$ when we compare equal volumes (i.e. $d=0$). The relative depth of the junction and GB can then be computed as follows
$\begin{split}\frac{d_{\rm J}}{d_{\rm GB}}&=\frac{H-z_{n}(\theta,R)}{H-z_{n}(\theta,a)}=\frac{H/a}{H/a-\tilde{z}_{n}(\theta,1)}\\ &=\frac{1}{1-\frac{\tan(\pi/n)}{2\tilde{V}_{\rm C}(\theta)}\tilde{z}_{n}(\theta,1)}.\end{split}$ (S-25)
When the film is initially flat, we expect at early times that the groove progresses almost ideally following the early prediction reported in Ref. [8], with the grain interior acting as a half-infinite space. Thus, the depth of the groove is approximately (for small $\theta$) $d_{\rm GB}(t)\approx 0.78\tan(\theta)(Bt)^{1/4}$ where $B$ is the rate constant for surface diffusion. When $r^{\rm 3D}\gg r_{\rm c}^{\rm 3D}$, rupture occurs at early times, where this can be considered a good approximation. Assuming that rupture initiates at the junction, as evidenced also in the main text, and that the relative depth ratio derived approximates also dynamic junctions well, Eq. (S-25) can be used to estimate the critical rupture time, thus reading $t_{\rm C}=H^{4}/B(0.78\tan(\theta)d_{\rm J}/d_{\rm GB})^{-4}$. When $r^{\rm 3D}\gtrsim r_{\rm c}^{\rm 3D}$ rupture may occur later than $t_{\rm C}$ as the finite grain size slows down the grooving dynamics.